# Polarized $W^{+}W^{-}$ pairs at the LHC: Effects from bottom-quark induced processes at NLO QCD+EW

### v1

@@ 5 Conclusions
In this work, we have calculated the bottom-quark induced contribution to the doubly polarized cross sections of $W^{+}W^{-}$ pair production at the LHC using a fully leptonic decay mode at NLO QCD+EW level.
After the subtraction of the on-shell $tW$ production, this contribution is very small ($1\%$ for the YesVeto setup, $2\%$ for NoVeto) for the integrated unpolarized cross section, in agreement with the result of Ref. Baglio:2013toa . However, this effect is not equal on the four polarized cross sections. While being negligible for the TT polarization, it is largest for the LL case, at $9\%$ ($13\%$) for the YesVeto (NoVeto) setup. Beyond the integrated results, differential cross sections for individual polarizations have been presented and discussed, showing a distinct shape of the bottom-quark induced contribution to the LL polarization in some distributions. The magnitude of this effect can be large, e.g. reaching $30\%$ at $p_{T,W}\approx 100$ GeV. These results show that the LL polarization is more sensitive to the third-generation quark contribution.
@@
In addition, we have computed a bound of the $tW$ interference and found that it can reach $-4.4\%$ ($+1.0\%$) for the integrated LL cross section for the YesVeto (NoVeto) setup. For the kinematic distributions presented in this paper, this interference effect is negligible for NLO EW corrections, being less than $1\%$ in absolute value. For NLO QCD case, it can however be very large at some phase space regions, reaching $-20\%$ level for the LL polarization. Remarkably, we found that the $tW$ interference effect is significantly smaller for the NoVeto setup in comparison to the YesVeto one, for almost all differential distributions, except for the case of $\cos(\theta_{e}^{WW})$ angular distribution where the interference is large at the edges for both cut setups. One should be careful about this effect when subtracting the $tW$ contribution in precision measurements.

### v2

@@ 5 Conclusions
In this work, we have calculated the bottom-quark induced contribution to the doubly polarized cross sections of $W^{+}W^{-}$ pair production at the LHC using a fully leptonic decay mode at NLO QCD+EW level in the five-flavor scheme.
After the subtraction of the on-shell $tW$ production, this contribution is very small ($1\%$ for the YesVeto setup, $2\%$ for NoVeto) for the integrated unpolarized cross section, in agreement with the result of Baglio:2013toa ; Gehrmann:2014fva . However, this effect is not equal on the four polarized cross sections. While being negligible for the TT polarization, it is largest for the LL case, at $9\%$ ($13\%$) for the YesVeto (NoVeto) setup. Beyond the integrated results, differential cross sections for individual polarizations have been presented and discussed, showing a distinct shape of the bottom-quark induced contribution to the LL polarization in some distributions. The magnitude of this effect can be large, e.g. reaching $30\%$ at $p_{T,W}\approx 100$ GeV. These results show that the LL polarization is more sensitive to the third-generation quark contribution.
@@
In addition, we have computed a bound of the $tW$ interference and found that it can reach $-4.4\%$ ($+1.0\%$) for the integrated LL cross section for the YesVeto (NoVeto) setup. For the kinematic distributions presented in this paper, this interference effect is negligible for NLO EW corrections, being less than $1\%$ in absolute value. For NLO QCD case, it can however be very large at some phase space regions, reaching $-20\%$ level for the LL polarization. Remarkably, we found that the $tW$ interference effect is significantly smaller for the NoVeto setup in comparison to the YesVeto one, for almost all differential distributions, except for the case of $\cos(\theta_{e}^{WW})$ angular distribution where the interference is large at the edges for both cut setups.
Scale uncertainties on the signal and on the $tW$-interference bound have been provided using the direct scale variation method. For the $tW$-interference bound, we found that the width of the scale-uncertainty band increases with the magnitude of the bound. Except for the edges of the $\cos(\theta_{e}^{WW})$ distribution, the uncertainty band covers the case of vanishing $tW$-interference. In addition, at the $|\cos(\theta_{e}^{WW})|=1$ edges, the scale uncertainties on the signal are large, overwhelming the interference effect. It therefore remains hard to decide on the importance of the $tW$ interference for polarization measurements.